Circular No. 5787

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 A. V. Filippenko and T. Matheson, University of California at Berke-

ley, communicate: “Spectra (range 300-1000 nm, resolution 0.5 nm) of SN 1993J obtained on May 11 UT with the Shane 3-m reflector at Lick Observatory, in comparison with previous spectra, reveal that SN 1993J is gradually transforming itself into a type-Ib supernova. Prominent He I P-Cyg profiles have emerged, and the strength of Hα emission is decreasing. The ‘double peak’ appearance of Hα (cf. IAUC 5776, 5779), which has grown even more striking, is produced by Hα blended with He I 667.8 nm. Previously, the only other supernova known to have clearly experienced a metamorphosis from type II to type Ib (or Ic) is SN 1987K (Filippenko 1988, A.J. 96, 1941), although the possible development of He I lines could not be monitored due to the object’s proximity to the sun. We predict that after a few months, the optical spectrum of SN 1993J will be dominated by strong, broad [O I], [Ca II], and Ca II emission lines, with Hα very weak or absent. Thus, SN 1993J might be called a ‘type IIb’ supernova, following the suggestion of Woosley et al. (1987, Ap.J. 318, 664) that such objects could exist. The initial mass of the progenitor was probably 10-20 M_{⊙}, but mass transfer onto a close companion led to the removal of all but a small amount (few tenths of a solar mass) of the hydrogen envelope prior to the explosion. This interpretation is supported by preliminary theoretical studies of the bolometric light curve being conducted by S. E. Woosley and his collaborators (private communication).”

PSR 1823–13 J. P. Finley and H. Ogelman, University of Wisconsin at Madison,

report: “We have detected the 101-ms pulsar PSR 1823–13 (dynamic age ∼ 22 000 yr) in the x-ray band 0.1-2.4 keV during an 8700-s observation carried out between 1992 Oct. 12 and 14 with the PSPC aboard ROSAT. The observed background-subtracted count rate, including vignetting and deadtime corrections, is 0.006 ± 0.001 count/s. The source counts reside above 0.7 keV, implying a hard, absorbed x-ray spectrum. No pulsations were detected, and we derive a 90-percent confidence level upper limit of 35 percent on the pulsed fraction under the assumption of a sinusoidal modulation at the radio period.”

1993 May 11 Daniel W. E. Green